OpenStudy (anonymous):

what is the distance between the points (-2,8)and(-10,2) i said a 100units is that correct

OpenStudy (anonymous):

d=sqrt[(-10+2)^2+(2-8)^2] d=sqrt(64+36) d=sqrt100 d=10